OpenStudy (anonymous):

2log(3-x)=log2+log(22-2x)

OpenStudy (anonymous):

log (44-4x) = log (3-x)^2

OpenStudy (aravindg):

which base?

OpenStudy (anonymous):

therefore, 9 - 6x + x^2 = 44 - 4x. so x^2 - 2x - 35 = 0 therefore x = 7, -5

OpenStudy (anonymous):

just log means base 10.

OpenStudy (anonymous):

its to the base 10

OpenStudy (anonymous):

thanx a lot tanvi. thats really helpful.

OpenStudy (aravindg):

-5

OpenStudy (aravindg):

no just log means base e

OpenStudy (anonymous):

no problem :) no, see on your calci. just log is base 10, ln is base e.

OpenStudy (anonymous):

no in canada it means to base 10

OpenStudy (anonymous):

ohhhh, canada ;)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

where are you from?

OpenStudy (anonymous):

india :)

OpenStudy (anonymous):

thats awesome.... visited india lots

OpenStudy (anonymous):

cool :)